The product of 6 and nine times a number is the same as the difference of ten times the number and 2. What is the number?
Answer by checkley77(12844)

6*9x=10x-2
54x=10-2
54x-10x=-2
44x=-2
x=-2/44
x=-1/22 ans.
Proof:
54*-1/22=10*-1/22-2
-54/22=-10/22-2
-54/22=(-10-2*22)/22
-54/22=(-10-44)/22
-54/22=-54/22